two glass plates are touching at one end and separated by a thin wire at the other end. When a monochromatic parallel beam of wavelength 4200A incident normally on the glass plates is reflected, an interference pattern of 30 fringes is observed. If the wavelength of light used is taken 7000A instead of 4200A, the number of fringes observed will be?
A = Angstrom 

Grade:12th pass

1 Answers

Vishal
one year ago
Thetha is angle between the two glass plate 
 
B=λ/2θ
Number of fringes*fringe length=total length
30*λ/2 θ=x*λ/2 θ
30*4200=x*7000
X=18